OpenStudy (anonymous):

. Use the definition of the derivative to show that for the function f(x)=3/1x^2−5, the derivative is f(x)=3/2x.

OpenStudy (anonymous):

lim h-->0 f(x+h) - f(x) / h

OpenStudy (anonymous):

plug and chug

OpenStudy (anonymous):

I dont understand this what so ever

OpenStudy (anonymous):

you make one big equation...anywhere you see x you replace it with x+h

OpenStudy (anonymous):

then you subtract that from the original equation...and most of it should cancel...then you put 0 in for h and there you go

OpenStudy (anonymous):

when i plug it in will i get [1/3(x+h)-5-(1/3x ^{2}-5)\]

OpenStudy (anonymous):

isn't it 3/x^2-5?

OpenStudy (anonymous):

limit the h to 0 and you'll get what you need f(x) = x^n; f'(x) = nx^n-1 f(x) = constant; f'(x) = 0

OpenStudy (anonymous):

tx...he has to use the definition of a derivative

OpenStudy (anonymous):

no srry its 1/3

OpenStudy (anonymous):

so ti's 1/3x-5?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

1/3x^2-5?

OpenStudy (anonymous):

okay so then your gonna get (1/3(x+h)^2 -5 ) - (1/3x^2-5)

OpenStudy (anonymous):

foil out the (x+h)^2

OpenStudy (anonymous):

so you get (1/3(x^2+2h+h^2)-5)- (1/3x^2-5)

OpenStudy (anonymous):

(1/3x^2+1/3h^2-5)-(1/3x^2-5)

OpenStudy (anonymous):

now distribute the - sign for the 2nd equation

OpenStudy (anonymous):

where did 2h come from

OpenStudy (anonymous):

(x+h)(x+h) = x^2 +2xh + h^2

OpenStudy (anonymous):

thats where im messin up at im just squaring them and now that you said foil im understanding it. imma try and wrk it out

OpenStudy (anonymous):

yea....you'll see after you distribute the - sign...alot will cancel

OpenStudy (anonymous):

\[\frac{\frac{1}{3}(x+h)^2-5-(\frac{1}{3}x^2-5)}{h}\] \[=\frac{\frac{1}{3}(x^2+2xh+h^2)-5-\frac{1}{x}x^2+5}{h}\] \[=\frac{\frac{1}{3}(2xh+h^2)}{h}\] \[=\frac{\frac{1}{3}h(2x+h)}{h}\] \[=\frac{1}{3}(2x+h)\] now let h go to 0 get your answer

OpenStudy (anonymous):

so wouldnt b -2/3x